# BurstGPT: A Real-World Workload Dataset to Optimize LLM Serving Systems

### v3

@@ 4.1 Implementation
Figure 12 presents an overview of the workload generator. The system comprises four main components: the Prompt Sampler, the Prompt Pool, the Concurrency Generator, and the HTTP client. The HTTP client is included using the efficient and asynchronous aiohttp framework, a strategic choice aimed at optimizing interactions with the LLM inference server. The prompt pool can be constructed from any pre-processed prompt dataset. An index with token numbers as keys and prompt indices as values will be established in the Prompt Pool to mitigate the overhead of prompt sampler queries. In cases where multiple prompt indices correspond to a single token number, a prompt is randomly selected for return. Once the prompt sampler retrieves a prompt from the Prompt Pool to enable workload concurrency, the concurrency generator utilizes asyncio.sleep for timing the requests. This process concurrently cedes CPU control back to the prompt sampler, enabling it to proceed with the generation of the subsequent request.
@@
Requests of specified lengths are sampled from the Prompt Pool at stochastic intervals that are generated by Concurrency Generator in the Workload Generator. This concurrency pattern is treated as a time series, which inputs it into the Inference Engine via HTTP for performance evaluation purposes. The benchmark suite systematically adjusts RPS if we use the vanilla trace, or the parameters $\lambda$, $\alpha$, and $\beta$ according to a predetermined sequence, if we use the modeled trace, thereby generating average and instantaneous performance metrics for the user.
4.2 Scaling BurstGPT to Any Scale
@@
4.3 Matching Request-response Patterns with Real-world Dialogues
For request lengths, in Section 3 and related simulations [5], it is observed that the request length follows a Zipf distribution. The sampler samples request lengths using a variable parameter $\lambda$, influencing the frequency of shorter requests in input streams.
@@ 4.3 Matching Request-response Patterns with Real-world Dialogues
After determining request lengths, the suite initializes a CPU buffer for loading a truncated prompt dataset pool. This dataset can be sampled from existing GPT conversation dialogues, such as Alpaca [29]. The dataset truncation and request length sampling adhere to the predefined Zipf distribution. The buffer size is user-configurable to balance with the CPU’s memory capacity. Requests of specific lengths are then selected from the prompt dataset and sent to the server at intervals, where they queue for processing.
4.4 Metrics and Setups
In serving systems, Quality of Service (QoS) is often measured by latency and throughput [23]. Latency refers to the completion time of each token, while throughput is the number of tokens processed within a given period. In addition to efficiency, the stability and reliability of LLM serving are particularly important to users, yet these aspects are often overlooked. In this evaluation, we also focus on latency jitters to assess stability and failure rates to assess the reliability of LLM serving.
@@ 4.4 Metrics and Setups
Our evaluation is conducted on Llama-2-13b-chat on an A800 GPU server. The CPU is a 64-core Intel(R) Xeon(R) Platinum 8358P @ 2.60GHz, with 1 TB memory. In experiments that uses modeled scaling, we set the initial parameters $\alpha_{Gamma}$ and $\beta_{Gamma}$ of the Gamma distribution in BurstGPT to 0.5 and 2, respectively, for 30% GPU KV Cache utilization. The request length parameter $\theta_{zipf}$ is set to 1.1. We use requests in ShareGPT as the request pool.
Figure 13: The impact of concurrency of BurstGPT Conversation and BurstGPT API traces on the latency, throughput, KV cache usage and failure rates of vLLM serving on a single A800 GPU. Using the first 2000 queries for both traces, we scale BurstGPT by a factor of 10 to reach an RPS of 0.326. To reach the same RPS, MAF2 is scaled with a parameter of 1.234, ensuring both complete in 6126 seconds.
4.5 Evaluations on BurstGPT and MAF2 Workloads
@@
Figure 14: Average token latency and throughtput of vLLM serving Llama-2-13b-chat on an A800 GPU server with BurstGPT and MAF2 workloads.
@@
In this evaluation, we assess the efficiency and stability of vLLM [25] using BurstGPT and MAF2. We set the same RPS to 0.326 to scale both MAF2 and BurstGPT. Both experiments completed in 6126 seconds. Figure13 shows the serving system’s instantaneous metrics under the two workloads. The results indicate that the request burstiness of concurrency in BurstGPT is higher, whereas the request distribution of the MAF2 workload is more uniform. As we can see from Figure 14, the average latency and throughput are higher for BurstGPT compared to MAF2, indicating a larger system resource occupation by BurstGPT. In conclusion, although both methods share the same RPS, the more bursty concurrency in BurstGPT introduces more challenges to the serving system, particularly in terms of KV cache management, compared to MAF2.
4.6 Evaluations on Request Scheduling Policies
Using the aforementioned serving system setups, we randomly sample a time range on BurstGPT and build two subsets with conversation workloads and API workloads and then used Modeled Scaling to derive the corresponding parameter sequences for generating workloads. We use this two generated workloads separately evaluate three scheduling strategis i.e. fisrt-come-first-serve (FCFS), short-request-first (SRF), long-request-fisrt (LRF), on vLLM.
For FCFS, in Figure 15, the failure rate is low on conversation trace, whereas it turns high due to the extreme burstiness in API service. Then, as illustrated in the Figure 16, the efficiency of the various scheduling methods may shift across different types of real-world workloads. Comparing FCFS with SRF and LRF on average values in terms of efficiency and stability, we can find that the optimization on conversation service when LRF is superior to FCFS in latency and and latency jitters does not stand in API service. This interesting finding inspire us that even using real-world workloads, an optimization on one workload does not necessarily generalize to other workloads.
Figure 15: Impact of concurrency in conversation and API serving on the latency, throughput, KV cache utilization and failure rates of vLLM using FCFS. The evaluation is conducted with the Llama-2-13b-chat model on with modeled scaling of BurstGPT on an A800 GPU server.
@@ 4.6 Evaluations on Request Scheduling Policies
Figure 16: Average token latency and throughtput of FCFS, SRF and LRF on vLLM serving Llama-2-13b-chat on an A800 GPU server with BurstGPT Conversations and API workloads.
5 Limitations
The data collected is from a single service provider. Including more LLM service providers from different cloud services would help demonstrate the universality of BurstGPT. Additionally, the current trace would be enhanced by including characteristics such as session ID, the number of conversations in each session, and response time. We plan to improve BurstGPT in future versions by incorporating the above information to further enhance LLM serving optimizations. Also, due to space limits, the use case does not include workload provisioning. However, our analytical study demonstrates request-response patterns can be beneficial for future provisioning work.
6 Broader Impacts and Conclusions
@@
This research emphasizes incorporating real-world workload data to enhance LLM serving systems. It identifies a critical gap: the need for such data in optimizing these systems. In this work, we introduce the real-world LLM workload BurstGPT from Azure OpenAI GPT service. We provide a detailed analysis of workload patterns, including concurrency patterns in LLM services. These traces enable a deeper understanding of serving system performance under complex and variable workloads, highlighting challenges and opportunities in workload-aware LLM serving optimizations.
Our evaluations on BurstGPT integrate concurrency modeling with real-world request contents to comprehensively assess LLM serving systems. Our assessments reveal a previously unidentified performance degradation of these systems under real-world workloads.
@@
We encourage the use of BurstGPT in optimizing and evaluating serving systems to improve the efficiency, stability, and reliability of LLM services under realistic workload conditions. We also advocate for data-driven methodologies in the development of LLM systems and plan to collect and share more workload traces for extensive research.
7 Acknowledgement
@@
This work was partially supported by the National Natural Science Foundation of China under Grant No. 62272122, a Hong Kong RIF grant under Grant No. R6021-20, and Hong Kong CRF grants under Grant No. C2004-21G and C7004-22G. Also supported by High-performance Computing Platform of The Hong Kong University of Science and Technology (Guangzhou).

### v4

@@ 4.1. Implementation
Figure 12 presents an overview of the workload generator. The system comprises four main components: the Prompt Sampler, the Prompt Pool, the Concurrency Generator, and the HTTP client. The HTTP client is included using the efficient and asynchronous aiohttp framework. An index with token numbers as keys and prompt indices as values will be established in the Prompt Pool to mitigate the overhead of prompt sampler queries and a prompt is randomly selected for return. Once the prompt sampler retrieves a prompt from the Prompt Pool to enable workload concurrency, the concurrency generator times requests.
@@
Requests of specified lengths are sampled from the Prompt Pool at stochastic intervals that are generated by Concurrency Generator in the Workload Generator. This concurrency pattern is treated as a time series, which inputs it into the Inference Engine via HTTP for performance evaluation purposes. The benchmark suite systematically adjusts RPS if we use the vanilla trace, or the parameters $\lambda$, $\alpha$, and $\beta$ according to a predetermined sequence, if we use modeled traces, thereby generating average and instantaneous performance metrics for users.
Figure 13. The impact of concurrency of BurstGPT Conversation and BurstGPT API traces on the latency, throughput, KV cache usage, and failure rates of vLLM serving on a single A800 GPU. Using the first 2000 queries for both traces, we scale BurstGPT by a factor of 10 to reach an RPS of 0.326. To reach the same RPS, MAF2 is scaled with a parameter of 1.234, ensuring both are complete in 6126 seconds.
4.2. Scaling BurstGPT to Any Scale
@@
4.3. Matching BurstGPT with Dialogues
For request lengths, in Section 3 and related simulations (Li et al., 2023b), it is observed that the request length follows a Zipf distribution. The sampler samples request lengths using a variable parameter $\lambda$, influencing the frequency of shorter requests in input streams.
@@ 4.3. Matching BurstGPT with Dialogues
After determining request lengths, the suite initializes a CPU buffer for loading a truncated prompt dataset pool. This dataset can be sampled from existing GPT conversation dialogues, such as Alpaca (Li et al., 2023a). The dataset truncation and request length sampling adhere to the predefined Zipf distribution. The buffer size is user-configurable to balance with the CPU’s memory capacity. Requests of specific lengths are then selected from the prompt dataset and sent to the server at intervals, where they queue for processing.
4.4. Metrics and Setups
In serving systems, Quality of Service (QoS) is often measured by latency and throughput (Miao et al., 2023). Latency refers to the completion time of each token, while throughput is the number of tokens processed within a given period. In addition to efficiency, the stability and reliability of LLM serving are particularly important to users, yet these aspects are often overlooked. In this evaluation, we also focus on latency jitters to assess stability and failure rates to assess the reliability of LLM serving.
@@ 5. Demo Evaluations
Our evaluation is conducted on Llama-2-13b-chat on an A800 and A6000 GPU server. In experiments that use modeled scaling, we set the initial parameters $\alpha_{Gamma}$ and $\beta_{Gamma}$ of the Gamma distribution in BurstGPT to 0.5 and 2, respectively, for 30% GPU KV Cache utilization. The request length parameter $\theta_{zipf}$ is set to 1.1. We use requests in ShareGPT as the request pool.
5.1. Evaluation: BurstGPT v.s. MAF2
@@
Figure 14. Average token latency and throughtput of vLLM serving Llama-2-13b-chat on an A800 GPU server with BurstGPT and MAF2 workloads.
@@
In this evaluation, we assess the efficiency and stability of vLLM (Kwon et al., 2023a) using BurstGPT and MAF2. We set the same RPS to 0.326 to scale both MAF2 and BurstGPT. Both experiments were completed in 6126 seconds. Figure13 shows the serving system’s instantaneous metrics under the two workloads. The results indicate that the request burstiness of concurrency in BurstGPT is higher, whereas the request distribution of the MAF2 workload is more uniform. As we can see from Figure 14, the average latency and throughput are higher for BurstGPT compared to MAF2, indicating a larger system resource occupation by BurstGPT. In conclusion, although both methods share the same RPS, the more bursty concurrency in BurstGPT introduces more challenges to the serving system, particularly in terms of KV cache management, compared to MAF2.
5.2. Evaluation: Micro Workload Variations
Figure 15. Impact of parameters $\alpha$, $\beta$, and $\lambda$ on average and p90 latency, average throughput, and average failure rates in vLLM. The evaluation is conducted with Llama-2-7b-chat and Llama-2-13b-chat models using A6000 and A800 GPU Servers. Experiment Settings: Larger $\alpha$ or smaller $CV$ and $\beta$ increase system burstiness; higher $\lambda$ correlates with shorter request likelihood. Note: Due to space limits, only results from the A6000 server are shown.
In this experiment, we model the request lengths to a Zipf distribution, as in (Li et al., 2023b). The sampler samples the request lengths in Zipf distribution with a variable parameter $\lambda$. We investigate the impact of varying specific variables: $\alpha$, $\beta$, and $\lambda$. We utilized the Llama-2-7b-chat and Llama-2-13b-chat models in our testbed. The results demonstrate a correlation between the adjustments in these variables and the observed system behavior.
We observe that minor changes in the parameter $\alpha$ or the $CV$ can lead to sudden request failures, significantly undermining the system’s reliability. However, these changes have less effect on system performance metrics such as average latency and throughput. This is because such changes in $\alpha$ are temporary and behave as quadratic functions. In contrast, alterations in $\alpha$ in a linear decline lead to ongoing performance degradation, particularly noticeable during request failures. These findings underscore the value of our trace analysis and provide essential insights for improving system efficiency in similar computational environments.
5.3. Demo Use: Request Scheduler Selection
@@
Figure 16. Average token latency and throughput of FCFS, SRF, and LRF on vLLM serving Llama-2-13b-chat on an A800 GPU server with BurstGPT Conversations and API workloads.
We randomly sample a time range on BurstGPT and build two subsets with conversation workloads and API workloads and then used Modeled Scaling to derive the corresponding parameter sequences for generating workloads. We use these two generated workloads separately to evaluate three scheduling strategies i.e. fisrt-come-first-serve (FCFS), short-request-first (SRF), and long-request-fisrt (LRF), on vLLM.
In Figure 16, the efficiency of the various scheduling methods may shift across different types of real-world workloads. Comparing FCFS with SRF and LRF on average values in terms of efficiency and stability, we can find that the optimization on conversation service when LRF is superior to FCFS in latency and and latency jitters does not stand in API service. This interesting finding inspires us that even using real-world workloads, an optimization on one workload does not necessarily generalize to other workloads.
5.4. Demo Use: Workload Provisioning
To adjust the resources allocated to services in a timely manner, cloud systems can predict system load to proactively scale resources up or down. The workload provisioning task is a time series forecasting problem. Specifically, we select the request count per time bin and the average number of tokens per request as the prediction targets. To construct features for the prediction model, we use historical load data from both short-term and long-term perspectives. For each time point, we incorporate load values from the past 3 time points (lag window) to capture short-term trends, along with statistical features (e.g., mean, variance) computed over a rolling window of up to 60 minutes to reflect long-term patterns. As shown in 3.1, the day of the week and the hour of the day also impact the workload, so these two features are included. The prediction target is the actual load value for that time point. We employ XGBoost (Chen and Guestrin, 2016) as the prediction model due to its simplicity and effectiveness. We compute the normalized mean absolute error (NMAE) and normalized mean square error (NMSE) between the prediction and the ground truth. The results are shown in Figure 17.
As shown, even without careful tuning, the models can predict future system load easily and accurately. The predicted load trends generally align well with the actual load, indicating significant opportunities for fine-grained load prediction modeling in real-world inference services, which lead to more precise service delivery.
We also investigate the performance of the prediction algorithm under different time intervals. As shown in Figure 18, predicting long-term patterns (10 minutes) is generally easier than predicting short-term patterns (1 minute), with the former achieving a 58% lower NMSE compared to the latter. This suggests that when using load prediction for serving, it is essential to balance the granularity of scheduling with the accuracy of predictions.
Figure 17. Request count and mean request token number predictions. The figure only shows a random interval of 6 hours. The time bin interval is 1 minute. Left: Request count predictions, the NMAE is 0.73 and the NMSE is 0.48. Right: Mean request token number predictions, the NMAE is 0.72 and the NMSE is 0.67.
Figure 18. Request count predictions. The figure only shows a random interval of 6 hours. Left: Request count predictions of Conv. with time bin interval 1 minute, the NMAE is 0.66 and the NMSE is 0.50. Right: Request count predictions of Conv. with time bin interval 10 minutes, the NMAE is 0.32 and the NMSE is 0.21.
6. BurstGPT in Industry
We provide an example of how we use BurstGPT for industrial use in Huawei during our development of $PD$ disaggregation systems. As a common practice in the industry, given a fixed serving system scale, developers use goodput (Zhong et al., 2024) to evaluate the system’s capacity, i.e., the maximum request rate that can be served while meeting the SLO goal (e.g., 90%) for each GPU. Recently, disaggregating the prefilling and decoding processes during LLM inference has become the standard approach in LLM service deployment (Hu et al., 2024; Qin et al., 2024; Zhong et al., 2024). However, scheduling the disaggregation of prefilling and decoding can be challenging. The prefilling instances can be seen as the "Provider," which generates the first KV cache and token, while the decoding instances act as the "Consumer," accepting the KV cache and token to continue the auto-regressive generation. Thus, for a given system scale (i.e., the number of prefilling ($P$) and decoding ($D$) instances), the concurrency of requests, request lengths, and response lengths will determine the goodput of the system. As workloads vary rapidly in real-world serving, a fixed $PD$ ratio is not always ideal. Instead, a dynamic $PD$ ratio should be used.
Figure 20 and Figure 19 show examples of goodput at different PD ratios. Since we lack preliminary workload traces before the real-world deployment of our serving system, we use BurstGPT in our simulation platform to determine the optimal PD ratio under various user, model, and system workloads. The result is used to develop prototypes of the $PD$ disaggregation serving system at Huawei. In the simulation, we select the $P$ and $D$ instance ratio, then run simulations, gradually increasing QPS while monitoring the delay SLO achievement rate until the SLO is no longer met. The QPS at this point is considered the system’s capacity. We then fix the total number of $P$ and $D$ instances. First, considering a simple situation, we adjust the $PD$ ratio one time to fit the workload divergence in Figure 19, we observe that at 6:00, the $PD$ ratio switches from 2:6 to 6:2, improving the goodput of the serving system. This simulation serves as a solid prototype for real-world $PD$ disaggregation scheduling policies. Then, we develop a dynamic $PD$ ratio strategy. We use beam search to identify the best $P$$D$ ratio given the provisioned workload in BurstGPT. Instead of directly scaling the number of instances, we simulate scaling by adjusting the load size per instance, assuming a predefined instance launch time in seconds. As shown in Figure 20, with the provisioned workload, the system searches for the optimal $PD$ ratio to improve the goodput.
Figure 19. At 6:00, the PD ratio switches from 2:6 to 6:2, improving the goodput of the serving system.
Figure 20. Given 8 instances, comparison of different PD ratio and dynamic PD ratio based on workload provisioning and beam search during LLM serving.
These findings highlight the significant improvements in LLM serving achieved through data-driven modeling of user and model behavior. By carefully designing the scheduling algorithm, we can greatly enhance both system efficiency and QoS.
7. Broader Impacts and Conclusions
@@
This research emphasizes incorporating real-world workload data to enhance LLM serving systems. It identifies a critical gap: the need for such data in optimizing these systems. In this work, we introduce the real-world LLM workload BurstGPT from Azure OpenAI GPT service. The traces and evaluations reveal possible performance degradation of serving systems under real-world workloads, highlighting challenges and opportunities in workload-aware LLM serving optimizations.
@@
We encourage using BurstGPT in optimizing and evaluating serving systems to improve the efficiency, stability, and reliability of LLM services under realistic workloads. We also advocate for data-driven methodologies in developing LLM systems in the future.
8. Acknowledgement
@@
This work that involves all hardware use was done and open-sourced while Yuxin Wang was a research intern at the Hong Kong University of Science and Technology (Guangzhou). It was partially supported by the National Natural Science Foundation of China under Grant No. 62272122, a Hong Kong RIF grant under Grant No. R6021-20, and Hong Kong CRF grants under Grant No. C2004-21G and C7004-22G. Also supported by the the High-performance Computing Platform and the Red Bird MPhil Program of the Hong Kong University of Science and Technology (Guangzhou). For all closed-source models, evaluations are performed through their respective APIs.